OpenStudy (anonymous):

y=sqrt(-x+2)-1 (A)What was the basic function? (B) What transformation were applied? (C) Label final intercepts

OpenStudy (xapproachesinfinity):

\(\normalsize\color{green}{ so\ the\ parent\ function\ is\ f(x)=sqrt{x} }\)

OpenStudy (xapproachesinfinity):

\(\normalsize\color{green}{ \sqrt{x} }\)

OpenStudy (xapproachesinfinity):

\(\normalsize\color{green}{ if\ g(x)=f(x+1) \ it\ is\ shifted\ 1\ uniyt\ to\ the \ left\\ if \ g(x)=f(x-1)\ it\ is\ shifted\ 1 unit\ to\ the\ the\ right\\ if g(x)= f(x)+1 \ it\ is\ shifted\ 1\ unit\ up\\ if g(x)= f(x)-1 it\ is\ shifted \ down }\)

OpenStudy (xapproachesinfinity):

Can you apply this to that function?

OpenStudy (anonymous):

would the intercepts be x:(1,0) \[y:(0, \sqrt{2}-1)\]

OpenStudy (xapproachesinfinity):

if g(x)=-f(x) it reflected

OpenStudy (anonymous):

eh, is there something im not seeing?

OpenStudy (xapproachesinfinity):

y intercept is as you said. X intercepts you have to f(x)=0 so I don't know if you got the right answer

OpenStudy (anonymous):

ok can we start from the first one how would I find the basic function

OpenStudy (xapproachesinfinity):

No you're good

OpenStudy (xapproachesinfinity):

the basic function is rootx it is the parent function rootsquare

OpenStudy (anonymous):

so basic function is?\[\sqrt{x}\]

OpenStudy (anonymous):

thats what I got from your sentence

OpenStudy (xapproachesinfinity):

yes correct

OpenStudy (xapproachesinfinity):

for B you need to recognize what happens to that basic function that got us to the function above

OpenStudy (anonymous):

oh ok

OpenStudy (xapproachesinfinity):

I mean the transformation

OpenStudy (anonymous):

so it goes to the left 2 and down 1?

OpenStudy (xapproachesinfinity):

right!

OpenStudy (xapproachesinfinity):

and it is also reflected

OpenStudy (anonymous):

oh yea the x is negative

OpenStudy (xapproachesinfinity):

right!

OpenStudy (anonymous):

so then that's B, then C is the Final intercepts meaning the last x and y intercepts right?

OpenStudy (xapproachesinfinity):

Correct!

OpenStudy (anonymous):

which is what i put well thanks for the help!

OpenStudy (anonymous):

It gets easier the more I do it

OpenStudy (xapproachesinfinity):

Exactly!

OpenStudy (xapproachesinfinity):

You're welcome

OpenStudy (anonymous):

alright see ya

OpenStudy (xapproachesinfinity):

yes! it practice the more you do it the more it gets easier. the habit they say haha

OpenStudy (anonymous):

yea

OpenStudy (xapproachesinfinity):

Alright!